# Attention-Guided Perturbation for Unsupervised Image Anomaly Detection

### v1

@@ Abstract
Reconstruction-based methods have significantly advanced modern unsupervised anomaly detection. However, the strong capacity of neural networks often violates the underlying assumptions by reconstructing abnormal samples well. To alleviate this issue, we present a simple yet effective reconstruction framework named Attention-Guided Pertuation Network (AGPNet), which learns to add perturbation noise with an attention mask, for accurate unsupervised anomaly detection. Specifically, it consists of two branches, i.e., a plain reconstruction branch and an auxiliary attention-based perturbation branch. The reconstruction branch is simply a plain reconstruction network that learns to reconstruct normal samples, while the auxiliary branch aims to produce attention masks to guide the noise perturbation process for normal samples from easy to hard. By doing so, we are expecting to synthesize hard yet more informative anomalies for training, which enable the reconstruction branch to learn important inherent normal patterns both comprehensively and efficiently. Extensive experiments are conducted on three popular benchmarks covering MVTec-AD, VisA, and MVTec-3D, and show that our framework obtains leading anomaly detection performance under various setups including few-shot, one-class, and multi-class setups.
Index Terms:
Image Anomaly Detection, Reconstruction, Noise
I Introduction
@@
Automatically detecting and localizing anomalies in images, ie, visual anomaly detection (VAD), has extensive applications in manufacturing [1, 2]. However, this task typically encounters the challenge of cold-start [3] where gathering normal samples is easy, whereas acquiring anomalous samples is costly or infeasible, which prohibits supervised learning. Therefore, significant effort has been devoted to developing unsupervised anomaly detection (UAD) algorithms [4, 5, 6, 7, 8, 9] that learn to model normality distribution with solely normal images.
Figure 1: Comparisons between our and existing perturbation strategies in reconstruction paradigm for anomaly detection. Compared with (a) existing networks adopting fixed or random perturbations, (b) we introduce to learn to guide the perturbation with sample-aware masks from the network and help to learn a better reconstruction model.
@@
Among existing UAD works, reconstruction-based methods [10, 11] hold great promise and have been extensively studied. Generally, they assume that reconstruction networks exposed to normal-only training images can reconstruct normal ones well, but find difficulty in reconstructing abnormal ones. Thus, an anomaly map can be computed by analyzing the reconstruction error between the input and its reconstructed counterpart. As per this assumption, some methods attempt to reconstruct input images with popular generative models, eg, AutoEncoder [12, 13, 4], GANs [14, 15, 16], or Difusion-based models [17, 18]. However, reconstructing raw pixel values usually fails when normal and anomalous regions share similar values [19]. Another direction is to learn to reconstruct features since they provide more contextual information than raw pixels [19, 20, 5]. With the availability of general-purpose large pretrained networks, there has been a growing interest in learning a reconstruction network to reconstruct features from pretrained networks which have been pretrained on large-scale datasets, e.g., ImageNet, and provide discriminative representations.
@@
Despite the outstanding performance, it has been commonly observed [5, 19, 3] that reconstruction-based methods often encounter the “identity shortcut” issue, a phenomenon of reconstructing anomalies well, due to their uncontrollable generalizability. To alleviate this issue, many strategies [21, 22, 3] have been presented. One simple yet effective strategy is to add perturbations to the input to suppress the reconstruction of anomalies (Fig. 1(a)). For instance, AESc [23] introduces a stain perturbation model to corrupt the image with an ellipse shape of irregular edges. During training, they simply restore the corruption. SCADN [16] generates fixed multi-scale striped masks to erase a part of regions in the input image and learn to reconstruct the missing information. DRAEM [3] obtains promising results by recovering pseudo-anomaly disturbed normal images for representation. By perturbing the input, the reconstruction network is forced to learn invariant patterns for reconstruction. Evidently, an effective perturbation scheme is critical. However, prior works simply apply fixed or random masks to perturb input images indiscriminately, regardless of their specific contextual characteristics. In real-world industrial settings with diverse categories and varying shapes and sizes of foreground objects, such strategies easily induce the reconstruction model to overfit the irrelevant background, leading to suboptimal performance.
@@
Our framework has several desirable merits. First, it identifies anomalies at high accuracy under both one-class and multi-class settings. By perturbating more aggressively on the important regions, we are in fact attempting to synthesize hard anomalous locations for training, which helps to learn more discriminative and compact boundaries to spot the abnormal patterns. Besides, the attention masks are sample-aware, accommodating images from diverse categories that exhibit different anomaly types. Second, it is more efficient to train. Anomalies in industrial images usually exhibit diverse structures and textures scattered across any foreground locations. Some anomalies can be easily identified by reconstructing them based on proximal pixels with similar colors, while anomalies with irregular structures near the contour or local edge pixels pose a challenge for reconstruction. Masking these critical areas is more conducive and efficient to representation learning for reconstruction. Lastly, it is a concise and effective framework. The whole main branch is just constructed with a pretrained backbone followed by a simple decoder, and the auxiliary branch will be removed after training, incurring no computation cost for inference.
@@
  • We present a simple yet effective framework named AGPNet, which integrates a reconstruction network with a novel perturbation scheme leveraging attention mechanisms. It can effectively alleviate “identical shortcut” by applying more targeted perturbations to crucial areas, resulting in superior performance under both multi-class and one-class settings compared with existing state-of-the-art methods.
@@
II Related Work
Due to the uncertain and scarce nature of visual anomalies, a dominant direction in this area is to formulate this problem as an unsupervised paradigm, also known as unsupervised anomaly detection. It assumes that all the images in the training set are normal while the test set contains both normal and abnormal images. A great deal of work has been proposed, for more comprehensive surveys, see [24, 25]. Generally, the methods can be broadly divided into three groups: synthesizing-based methods, embedding-based methods, and reconstruction-based methods.
Synthesizing-based methods [26, 27, 10, 3] synthesize anomalies for training. For instance, DRAEM [3] attempts to blend predefined texture images with normal samples. CutPaste [26] simply cuts an image patch and randomly pastes it to another image location. These methods may heavily rely on the quality of the synthesized anomalies. However, due to the uncertainty and unpredictability of anomalies, it is hard and even impossible to synthesize all types of real anomalies with high fidelity.
Embedding-based methods [2, 28, 29, 30] typically work by first utilizing pretrained models to embed normal images and then employing statistical models to model the normal distributions. During testing, samples are regarded as being abnormal if they are far away from the normal distributions. For instance, PaDiM [2] models the embedded patch features with multivariate Gaussian distribution. PatchCore [28] exploits the memory bank to store nominal features. However, this group of methods generally demands more computational resources to store the normal embeddings for the training set. In addition, the computing process to identify anomalies during inference is also time-consuming.
Reconstruction-based methods [5, 31, 32, 22] hold the assumption that the model that is trained only on normal data can reconstruct normal samples while finding difficulty in reconstructing anomalous during inference. However, this assumption sometimes is not satisfied because of the strong and uncontrollable generalization of deep neural networks. Therefore, many strategies have been proposed to address this issue. Some methods try to add prior from the images [22, 20, 33] as guidance for reconstructions. For instance, P-Net [22] proposes to feed the structure features into the reconstruction network to guide the reconstruction process. Some methods [23, 34, 35] frame anomaly detection as the image restoration problem by corrupting the selected content in the images and then learning to restore the corruptions. several methods [36, 37] take the idea of image inpainting by first masking the image randomly and then learning to recover it. Compared with these strategies, our method attempts to add noise at both the feature level and image level according to attention masks. The attention masks, which are computed based on the prior attention from our pretrained model and learnable attention from the decoder, better reflect the local details and their importance. By perturbing these regions more aggressively, our reconstruction network can have a better understanding of the intrinsic and important local normal patterns, which help suppress the anomaly reconstructions effectively.
III Method
Problem Formulation. Following previous works [3, 5], we formulate visual anomaly detection (VAD) as a reconstruction problem. Given a training set $\mathcal{X}_{train}=\{I_{normal}^{i}\}_{i=1}^{N}$ comprising only normal samples, our main goal is to train a reconstruction model $f(\cdot)$ to accurately identify both pixel and image-level anomalies in a test set $\mathcal{X}_{test}=\{I^{j}\}_{j=1}^{M}$ containing both normal and abnormal samples. Note that the problem of VAD can be further divided into two settings: one-class and multi-class. The one-class setting involves images from the same semantic categories, while the multi-class setting deals with images from diverse semantic categories. The latter is more challenging due to the diverse distributions of different categories, making it difficult to learn compact representations.
Model Overview. Our framework AGPNet can deal with both one-class and multi-class settings simultaneously. As shown in Fig. 2, it consists of two branches. The main reconstruction branch aims to reconstruct the inputs, while the auxiliary branch aims to generate attention masks to guide the perturbation process for the inputs. With the attention mask as perturbation guidance, the reconstruction network tends to learn invariant normal patterns at those crucial locations. Below we describe each component in detail.
III-A Reconstruction Branch
@@
Feature Extractor. Given an input RGB image $I\in\mathbb{R}^{H\times W\times 3}$, we feed it into a pretrained network $\phi(\cdot)$ and only take a subset of $L$ layer outputs. Any off-the-shelf pre-trained ViT-like model, which relies on attention mechanisms, can readily serve as the feature extractor. In this paper, we follow previous work [38] and adopt DINO [39] as our backbone, which is trained via self-distillation without labels, has demonstrated its efficacy in generating superior representations for VAD. Note that DINO keeps the resolution fixed along the layer hierarchy, and all the feature maps share the same dimension after packing them into 3D feature maps. We denote $F_{l}\in\mathbb{R}^{H_{f}\times W_{f}\times C_{f}}$ as the $l$-th output, and $\sigma(\cdot)$ as the layer normalization function. The representations from specific layers will undergo layer normalization individually, followed by summation:
@@
Another important reason to utilize DINO [39] is that it is a vison-transformer-based architecture and its attention weights can be leveraged to generate the attention mask to guide the decoder training with minimal efforts. Therefore, in addition to the $L$ outputs, we propose to reuse the associated attention maps. denoted by $A_{l}\in\mathbb{R}^{H_{l}\times W_{l}}$, to provide the important prior cues for the image to guide the training of the decoder. The produced attention weights are taken out and will further be forwarded to the attention-guided perturbation branch for further integration.
Decoder. The decoder aims to decode the pretrained features of normal data in the training set. However, without proper design, it easily falls into the “identity shortcut” issue, especially under the multi-class setting. Recently UniAD [5] attempts to alleviate this issue with a learnable-query decoder that incorporates additional learnable queries at each transformer layer. This design incurs additional costs. In our framework, we simply adopt a plain vision-transformer-based decoder thanks to an auxiliary attention-based perturbation branch. During training, the decoder is guided to decode the attention-guided perturbed representations derived from the feature extractor, rather than solely relying on the normal representation. Consequently, the “identity shotcut” issue will be greatly alleviated. Furthermore, our decoder has a more lightweight structure with only four layers, demonstrating strong empirical performance during inference.
III-B Attention-Guided Perturbation.
@@
As discussed earlier, the attention weights from the feature extractor directly reflect the importance of each localization in the feature maps [40]. Therefore, based on the extracted subset of attention weights $\{A_{l}\}_{l=1}^{L}$, we compute the prior attention mask by $A_{\text{prior}}=\Phi_{\text{aggr}}(\{A_{l}\}_{l=1}^{L})$, where $\Phi_{aggr}$ denotes the aggregation operation over the attention maps. For simplicity, we only apply element-wise average pooling on the attention maps.
Second, the attention weights within the decoder are learned to aggregate important contextual cues for reconstruction. Hence, we further propose to utilize it for attention mask generation. However, these attention weights fluctuate rapidly with high variance, especially at the beginning of the training. For stabilization, we employ mean-distillation [41], which includes an exponential-moving-average (EMA) version of the decoder as the teacher.
@@
Attention-Guided Noise. For unsupervised anomaly detection, synthesized anomalies play a critical role in learning a compact and discriminative boundary to discern normal and abnormal samples. Numerous approaches have been proposed [3, 26] for synthesizing anomalies that closely resemble real ones. Concerning the impracticality of synthesizing all types of real anomalies, we argue that focusing on critical locations and perturbing them at the feature level could be a more efficient strategy. Besides, by perturbing those critical and informative locations, the models are forced to reason the relations across different localizations more comprehensively. To this end, we simply add Gaussian noise weighted by the attention mask in the normal feature space to synthesize the hard abnormal samples. Specifically, a noise tensor $\mathcal{E}\in\mathbb{R}^{H_{f}\times\ W_{f}\times C_{f}}$ is first generated with each entry simply drawn from an i.i.d Gaussian distribution $\mathcal{N}(\mu,\sigma)$. Then we add the noise tensor to the normal features $F_{\text{clean}}$ for perturbation based on attention mask $A_{\text{final}}$ as follows:
@@
$\alpha(t)=\beta\cdot(\frac{t}{T}(p-m)+m),$ (5)
where $p$ represents the maximum noise intensity, $m$ represents the minimum noise intensity, $T$ represents the maximun training epoch and $t$ represents the current training epoch.
In addition to introducing noise at the feature level, we also incorporate noise at the image level, guided by the attention mask $A_{\text{final}}$, resulting in improved performance. To achieve this, we first upsample $A_{\text{final}}$ to match the image dimensions, producing $A_{\text{img}}\in\mathbb{R}^{H\times W}$. Subsequently, we progressively increase the mask ratio for binarization from 0.6 to 1.0. Gaussian noise is also applied within the masked regions.
III-C Loss Function

### v2

@@ Abstract
Reconstruction-based methods have significantly advanced unsupervised image anomaly detection involving only normal training images. However, it has been proven that modern neural networks generally have a strong reconstruction capacity and often reconstruct both normal and abnormal samples well, thereby failing to spot anomaly regions by checking the reconstruction quality. To prevent well-reconstructed anomalies, one simple but effective strategy is to perturb normal samples and then map perturbed versions to normal ones. Yet it treats each spatial position equally, disregarding the fact that the foreground locations are inherently more important for reconstruction. Motivated by this, we present a simple yet effective reconstruction framework named Attention-Guided Perturbation Network (AGPNet), which learns to add perturbations guided with an attention mask during training. Specifically, it consists of two branches, i.e., a reconstruction branch and an auxiliary attention-based perturbation branch. The reconstruction branch learns to reconstruct normal samples, while the auxiliary one aims to produce attention masks to guide the noise perturbation process for normal samples. By doing so, we are expecting to synthesize hard yet more informative anomalies for training, which enable the reconstruction branch to learn important inherent normal patterns both comprehensively and efficiently. Extensive experiments are conducted on several popular benchmarks covering MVTec-AD, VisA, and MVTec-3D, and show that AGPNet obtains leading anomaly detection results under few-shot, one-class, and multi-class setups.
keywords:
Unsupervised Image Anomaly Detection , Reconstruction , Perturbation , Attention.
journal: Computer Vision and Image Understanding
[…]
left=1.5cm, right=1.5cm, top=2.5cm, bottom=2.5cm
1 Introduction
@@
Automatically detecting and localizing anomalies in images, i.e., visual anomaly detection (VAD), has extensive applications in manufacturing Bergmann et al. (2019); Defard et al. (2021). However, this task typically encounters the challenge of cold-start Zavrtanik et al. (2021) where gathering normal samples is easy, whereas acquiring anomalous samples is costly or infeasible, which prohibits supervised learning. Therefore, significant effort has been devoted to developing unsupervised anomaly detection (UAD) algorithms Bergmann et al. (2019); You et al. (2022); Bergmann et al. (2020); Zhang et al. (2023); Deng and Li (2022); Schlüter et al. (2022) that learn to model normality distribution with solely normal images during the training process. During inference, images that are far away from the normality distribution are regarded to be anomalous.
@@
Among existing UAD works, reconstruction-based methods Zavrtanik et al. (2022); Liang et al. (2023) hold great promise and have been extensively studied. Generally, they assume that reconstruction networks exposed to normal-only training images can reconstruct normal ones well, but find difficulty in reconstructing abnormal ones. Thus, an anomaly map can be computed by analyzing the reconstruction error between the input and its reconstructed counterpart. As per this assumption, some methods attempt to reconstruct input images with popular generative models, eg, AutoEncoder Gong et al. (2019); Nguyen et al. (2019); Bergmann et al. (2019), GANs Sabokrou et al. (2018); Kim et al. (2020); Yan et al. (2021), or Difusion-based models He et al. (2024); Zhang et al. (2023). However, reconstructing raw pixel values usually fails when normal and anomalous regions share similar values You et al. (2022). Another direction is reconstructing features since they provide more contextual information than raw pixels You et al. (2022); Shi et al. (2021); You et al. (2022). With the availability of general-purpose large pretrained networks, there has been a growing interest in learning a reconstruction network to reconstruct discriminative features from models pretrained on large-scale datasets, e.g., ImageNet Deng et al. (2009).
Figure 1: Comparisons between ours and existing perturbation strategies in reconstruction paradigm for anomaly detection. Compared with (a) existing networks adopting fixed or random perturbations, (b) we introduce to learn to guide the perturbation with sample-aware masks from the network and help to learn a better reconstruction model.
@@
Despite the outstanding performance, it has been commonly observed You et al. (2022); Zavrtanik et al. (2021); You et al. (2022) that reconstruction-based methods often encounter the “identity shortcut” issue, a phenomenon of reconstructing anomalies well, due to their uncontrollable generalizability. To alleviate this issue, many strategies Haselmann et al. (2018); Zhou et al. (2020); Zavrtanik et al. (2021) have been presented. One simple yet effective strategy is to perturb input to suppress the reconstruction of anomalies (Fig. 1(a)). For instance, AESc Collin and De Vleeschouwer (2021) introduces a stain perturbation model to corrupt the image with an ellipse shape of irregular edges. During training, they simply restore the corruption. SCADN Yan et al. (2021) generates fixed multi-scale striped masks to erase a part of regions in the input image and learn to reconstruct the missing information. DRAEM Zavrtanik et al. (2021) obtains promising results by recovering pseudo-anomaly disturbed normal images for representation. By perturbing the input, the reconstruction network is forced to learn invariant patterns for reconstruction. Evidently, an effective perturbation scheme is critical. However, prior works simply apply fixed or random masks to perturb input images indiscriminately, regardless of their specific contextual and semantic characteristics. In real-world industrial settings with diverse categories and varying shapes and sizes of foreground objects, such strategies easily induce the reconstruction model to overfit the irrelevant background, leading to suboptimal performance.
@@
Our framework has several desirable merits. First, it identifies anomalies at high accuracy under both one-class and multi-class settings. By perturbating more aggressively on the important regions, we are in fact attempting to synthesize hard anomalous locations for training, which helps to learn more discriminative and compact boundaries to spot the abnormal patterns. Besides, the attention masks are sample-aware, accommodating images from diverse categories that exhibit different anomaly types. Second, it is more efficient to train. Anomalies in industrial images usually exhibit diverse structures and textures scattered across any foreground locations. Some anomalies can be easily identified by reconstructing them based on proximal pixels with similar colors, while anomalies with irregular structures near the contour or local edge pixels pose a challenge for reconstruction. Masking these critical areas is more conducive and efficient to representation learning for reconstruction. Lastly, it is a concise and effective framework. The whole main branch is just constructed with a pretrained backbone followed by a simple ViT-based decoder, and the auxiliary branch will be removed after training, incurring no computation cost for inference.
@@
Our contributions are summarized as follows:
  • 1.
    We present a simple yet effective framework named AGPNet, which integrates a simple reconstruction network with a novel perturbation scheme. Unlike existing perturbation schemes that treat each spatial location equally, it can apply more targeted perturbations to crucial areas for each sample to enhance the reconstruction process.
  • 2.
@@
  • 3.
    Extensive experiments show that AGPNet obtains superior performance on both multi-class and one-class settings, compared with existing state-of-the-arts. Besides, it can also be versatilely and effectively extended to the few-shot setup.
2 Related Work
2.1 Unsupervised Visual Anomaly Detection
Due to the uncertain and scarce nature of visual anomalies, a dominant direction in this area is to formulate this problem as an unsupervised paradigm, also known as unsupervised anomaly detection. It assumes that all images in the training set are normal while the test set contains both normal and abnormal images. A great deal of work has been proposed, for more comprehensive surveys, see Liu et al. (2024); Zipfel et al. (2023). Generally, they can be broadly divided into three groups: synthesizing-based, embedding-based, and reconstruction-based methods.
Synthesizing-based methods Li et al. (2021); Schlüter et al. (2022); Zavrtanik et al. (2022, 2021) synthesize anomalies for training. For instance, DRAEM Zavrtanik et al. (2021) attempts to blend predefined texture images with normal samples. CutPaste Li et al. (2021) simply cuts an image patch and randomly pastes it to another image location. These methods heavily rely on the synthesized anomaly quality. However, due to the uncertainty and unpredictability of anomalies, it is hard and even impossible to synthesize all types of real anomalies with high fidelity.
Embedding-based methods Defard et al. (2021); Roth et al. (2022); Liu et al. (2023); Bae et al. (2023) typically work by first utilizing pretrained models to embed normal images and then employing statistical models to model the normal distributions. During testing, samples are regarded as being abnormal if they are far away from the normal distributions. For instance, PaDiM Defard et al. (2021) models embedded patch features with multivariate Gaussian distribution. PatchCore Roth et al. (2022) exploits memory banks to store nominal features. However, they generally demand more computational resources to store the normal embeddings for the training set. In addition, the computing process to identify anomalies during inference is also time-consuming.
Reconstruction-based methods Zavrtanik et al. (2021); Schlüter et al. (2022); Zhou et al. (2020) hold the assumption that the model pretrained only on normal data can reconstruct normal samples while finding difficulty in reconstructing anomalous during inference. However, this assumption sometimes is not satisfied because of the strong and uncontrollable generalization of deep neural networks. Therefore, many strategies have been proposed to address this issue. Some methods try to add prior from the images Zhou et al. (2020); Shi et al. (2021); Xia et al. (2020) as guidance for reconstructions. For instance, P-Net Zhou et al. (2020) proposes to feed the structure features into the reconstruction network. Other methods Collin and De Vleeschouwer (2021); Ye et al. (2020); Liu et al. (2023) frame anomaly detection as the image restoration problem by corrupting the selected content in the images and then learning to restore the corruptions. several methods Li et al. (2020); Zavrtanik et al. (2021) take the idea of image inpainting by first masking the image randomly and then learning to recover it.
Despite promising results, the above methods focus on the “one for one” paradigm, which trains one separate model for each category and thus is more computationally expensive. Recently, some researchers have shifted their attention to another challenging “one for all” paradigm, which only has to train a single for all categories. UniAD You et al. (2022) is the pioneering work for this task. It observses that under this setup the reconstruction network more easily falls into “identical shortcut” problem. They further introduce learnable queries to the transformer network. Later, OmniAL Zhao (2023) proposes a panel-guided anomaly synthesis method to prevent the model from performing identical reconstruction. RLR He et al. (2024a) introduces learnable reference representations to force the framework to learn normal feature patterns explicitly for all categories. MambaAD He et al. (2024b) is the first to integrate the recent Mamba Gu and Dao (2023) architecture into reconstruction-based anomaly detection frameworks, which achieves efficient global contextual modeling while preserving local features.
Following previous works, our work is also reconstruction-based but deals with both “one for one” and “one for all” paradigms. Unlike most existing methods of designing sophisticated architectures, we present a perturbation scheme that could effectively force a simple plain ViT-based reconstruction decoder to learn invariant normal features. Compared with other perturbation schemes treating each spatial location equally, we use attention masks calculated from the model itself to guide the perturbation. By doing so, our reconstruction network can have a better understanding of the intrinsic and important local normal patterns, which help suppress the anomaly reconstructions effectively.
2.2 Few-shot Visual Anomaly Detection
Recent research has begun to focus on a more challenging task setting: few-shot anomaly detection. It deals with scenarios where only a few training samples are available. RegAD Huang et al. (2022) employs a siamese network architecture to train a registration network, calculating the Mahalanobis distance between the features of the test image and those of a support set containing only normal samples for anomaly detection. With the rise of vision-language models, WinCLIP+ Jeong et al. (2023) introduces CLIP Radford et al. (2021) for few-shot anomaly detection. PromptAD Li et al. (2024) further develops on CLIP, leveraging manually designed and learned anomaly prompts to expand the distance between anomalous and normal features, thereby assisting in anomaly detection.
Our framework can also be effectively extended to few-shot anomaly detection. Compared with existing works, our works does not rely additional memory bank or complex data augmentations. With the simple yet effective attention-based perturbation scheme, our framework is also sample-efficient, obtaining comparable performance to the current leading few-shot anomaly detection methods.
3 Method
Problem Formulation. Following previous works Zavrtanik et al. (2021); You et al. (2022), we formulate visual anomaly detection (VAD) as a reconstruction problem. Given a training set $\mathcal{X}_{train}=\{I_{normal}^{i}\}_{i=1}^{N}$ comprising only normal samples, our main goal is to train a reconstruction model $f(\cdot)$ to accurately identify both pixel and image-level anomalies in a test set $\mathcal{X}_{test}=\{I^{j}\}_{j=1}^{M}$ containing both normal and abnormal samples. Note that the problem of VAD can be further divided into two settings: one-class and multi-class. The one-class setting involves training individual models with images from each semantic category, while the multi-class setting aims to train a unified model with all images from diverse semantic categories. The latter is more computationally efficient due to the $\mathcal{O}(1)$ task complexity regardless of the semantic category number. Yet it is more challenging due to the diverse distributions of different categories, making it difficult to learn compact representations. Lastly, in addition to conventional unsupervised anomaly detection, our framework can also be versatilly extended to few-shot anomaly detection. In this setup, only a few images are provided to the framework for training.
3.1 AGPNet
3.1.1 Model Overview
As shown in Fig. 2, our framework mainly consists of two branches. The main reconstruction branch aims to reconstruct the inputs, while the auxiliary branch aims to generate attention masks to guide the perturbation process for the inputs. With the attention mask as perturbation guidance, the reconstruction network tends to learn invariant normal patterns at those crucial locations. After training, the framework can be used for anomaly detection directly by comparing the input and output of the decoder directly, with the discrepancy indicating the anomaly locations. Below we describe each component in detail.
@@ 3.1.1 Model Overview
3.1.2 Reconstruction Branch
@@
Feature Extractor. Given an input RGB image $I\in\mathbb{R}^{H\times W\times 3}$, we feed it into a pretrained network $\phi(\cdot)$ and only take a subset of $L$ layer outputs. Any off-the-shelf pre-trained ViT-like model, which relies on attention mechanisms, can readily serve as the feature extractor. In this paper, we follow previous work Reiss et al. (2022) and adopt DINO Caron et al. (2021) as our backbone, which is trained via self-distillation without labels, has demonstrated its efficacy in generating superior representations for VAD. Note that DINO keeps the resolution fixed along the layer hierarchy, and all the feature maps share the same dimension after packing them into 3D feature maps. We denote $F_{l}\in\mathbb{R}^{H_{f}\times W_{f}\times C_{f}}$ as the $l$-th output, and $\sigma(\cdot)$ as the layer normalization function. The representations from specific layers will undergo layer normalization individually, followed by summation:
@@
Another important reason to utilize DINO Caron et al. (2021) is that it is a vison-transformer-based architecture and its attention weights can be leveraged to generate the attention mask to guide the decoder training with minimal efforts. Therefore, in addition to the $L$ outputs, we propose to reuse the associated attention maps. denoted by $A_{l}\in\mathbb{R}^{H_{l}\times W_{l}}$, to provide the important prior cues for the image to guide the training of the decoder. The produced attention weights are taken out and will further be forwarded to the attention-guided perturbation branch for further integration.
Decoder. The decoder aims to decode the pretrained features of normal data in the training set. However, without proper design, it easily falls into the “identity shortcut” issue, especially under the multi-class setting. Recently UniAD You et al. (2022) attempts to alleviate this issue with a learnable-query decoder that incorporates additional learnable queries at each transformer layer. This design incurs additional costs. In our framework, we simply adopt a plain vision-transformer-based decoder thanks to an auxiliary attention-based perturbation branch. During training, the decoder is guided to decode the attention-guided perturbed representations derived from the feature extractor, rather than solely relying on the normal representation. Consequently, the “identity shotcut” issue will be greatly alleviated. Furthermore, our decoder has a more lightweight structure with only four layers, demonstrating strong empirical performance during inference.
3.1.3 Attention-Guided Perturbation.
@@
As discussed earlier, the attention weights from the feature extractor directly reflect the importance of each localization in the feature maps Zeiler and Fergus (2014). Therefore, based on the extracted subset of attention weights $\{A_{l}\}_{l=1}^{L}$, we compute the prior attention mask by $A_{\text{prior}}=\Phi_{\text{aggr}}(\{A_{l}\}_{l=1}^{L})$, where $\Phi_{aggr}$ denotes the aggregation operation over the attention maps. For simplicity, we only apply element-wise average pooling on the attention maps.
Second, the attention weights within the decoder are learned to aggregate important contextual cues for reconstruction. Hence, we further propose to utilize it for attention mask generation. However, these attention weights fluctuate rapidly with high variance, especially at the beginning of the training. For stabilization, we employ mean-distillation Tarvainen and Valpola (2018), which includes an exponential-moving-average (EMA) version of the decoder as the teacher.
@@
Attention-Guided Noise. For unsupervised anomaly detection, synthesized anomalies play a critical role in learning a compact and discriminative boundary to discern normal and abnormal samples. Numerous approaches have been proposed Zavrtanik et al. (2021); Li et al. (2021) for synthesizing anomalies that closely resemble real ones. Concerning the impracticality of synthesizing all types of real anomalies, we argue that focusing on critical locations and perturbing them at the feature level could be a more efficient strategy. Besides, by perturbing those critical and informative locations, the models are forced to reason the relations across different localizations more comprehensively. To this end, we simply add Gaussian noise weighted by the attention mask in the normal feature space to synthesize the hard abnormal samples. Specifically, a noise tensor $\mathcal{E}\in\mathbb{R}^{H_{f}\times\ W_{f}\times C_{f}}$ is first generated with each entry simply drawn from an i.i.d Gaussian distribution $\mathcal{N}(\mu,\sigma)$. Then we add the noise tensor to the normal features $F_{\text{clean}}$ for perturbation based on attention mask $A_{\text{final}}$ as follows:
@@
$\alpha(t)=\gamma\cdot(\frac{t}{T}(p-m)+m),$ (5)
where $\gamma$ represents the basic noise factor, $p$ represents the maximum noise intensity, $m$ represents the minimum noise intensity, $T$ represents the maximun training epoch and $t$ represents the current training epoch.
In addition to introducing noise at the feature level, we also incorporate noise at the image level, guided by the attention mask $A_{\text{final}}$, resulting in improved performance. To achieve this, we first upsample $A_{\text{final}}$ to match the image dimensions, producing $A_{\text{img}}\in\mathbb{R}^{H\times W}$. Subsequently, we progressively increase the mask ratio for binarization. Gaussian noise is also applied within the masked regions. Gradually increasing the mask ratio increases the difficulty of denoising the decoder, which helps model training.
3.2 Loss Function